Kayleybro2525:

The GCF of 12 and 15 is _____. Help you smart people This is in Math so help

Kayleybro2525:

come on

Timmyspu:

it is 3

Kayleybro2525:

how can you explane to me

Timmyspu:

sure

Timmyspu:

3 times 4 is 12 and 5 times 3 is 15 and there is no bigger number that goes into 12 and 15

Kayleybro2525:

can i text you and you are right

Timmyspu:

ya

jhonyy9:

yes @Timmyspu good job and explication

Timmyspu:

thx

jhonyy9:

12 = 3*4 15 = 3*5 what is the greatest common factor of these ?

jhonyy9:

@Kayleybro2525 do you understand it now ?